# A longitudinal Bayesian framework for estimating causal dose-response relationships

### v2

@@ Abstract
Existing causal methods for time-varying exposure and time-varying confounding focus on estimating the average causal effect of a time-varying binary treatment on an end-of-study outcome. Methods for estimating the effects of a time-varying continuous exposure at any dose level on the outcome are limited. We introduce a scalable, non-parametric Bayesian framework for estimating longitudinal causal dose–response relationships with repeated measures. We incorporate the generalized propensity score either as a covariate or through inverse-probability weighting, formulating two Bayesian dose–response estimators. The proposed approach embeds a double non-parametric generalized Bayesian bootstrap which enables a flexible Dirichlet process specification within a generalized estimating equations structure, capturing temporal correlation while making minimal assumptions about the functional form of the continuous exposure. We applied our proposed approach to a motivating study of monthly metro-ridership data and COVID-19 case counts from major international cities, identifying causal relationships and the dynamic dose–response patterns between higher ridership and increased case counts.
Keywords: Dose-response relationship; Generalized Bayesian bootstrap; Dirichlet processes; Longitudinal data analysis; COVID-19; Transportation engineering;
@@ 1 Introduction
Figure 1: Longitudinal causal DAG between the mass of metro ridership (treatment variable, $D$), confounders ($X$) and the number of COVID-19 cases ($Y$) from time $t$ to $t+1$.
$Y_{t}$$D_{t+1}$$X_{t+1}$$X_{t}$$Y_{t+1}$$D_{t}$}
@@
Using the standard causal inference notation, we define our problem in relation to the triple $(Y,D,X)$, indexed by $t$ to denote time: where $Y$ is the outcome (COVID-19 cases in the city), $D$ is the treatment (urban mass transit ridership), $X$ is a set of time-varying confounders We are interested in the causal effects of $D$ on $Y$. Figure 1 shows a directed acyclic graph (DAG) of our assumed time-varying causal relationships between mass transit ridership and disease propagation at the city level. The figure shows an autoregressive structure in $Y$, along with contemporaneous and lagged structure in both treatment effect and confounding. Therefore, we face two key challenges in estimating this causal effect. First, we must adjust for the existence of confounding influences, $X$ and $U$, which determine both transit ridership and disease propagation. These include time-varying (i.e., $X$) factors such as the severity of containment measures on human mobility (e.g. lockdown measures), urban mobility characteristics, and vaccination rates. Second, there are dynamic effects presented such that past outcomes act as contemporaneous confounders. Thus, we observe a repeatedly measured outcome, a time-dependent treatment and a set of dynamic time-varying and time-invariant confounders. For estimation, we have monthly (longitudinal) city level data available which measure Covid cases, mass transit ridership, mobility containment measures, and a set of mobility characteristics for the period between 1 March 2020 and 23 January 2022. Longitudinal data provide the temporal resolution needed to understand how public transit ridership and COVID-19 incidence evolve and intersect. To determine causality between the repeatedly measured COVID-19 cases and public transportation ridership, we thus require a robust model that explicitly builds the temporal dependencies illustrated in Figure 1.
@@
Most existing methods for causal estimation of a continuous-exposure are developed for the point-treatment framework, including the generalized propensity score-based (GPS) approaches (Robins et al., 2000; Hirano and Imbens, 2004; Galvao and Wang, 2015), g-computation and outcome regression-based approaches (Hill, 2011), and a doubly robust estimator considering combining both treatment and outcome models (Kennedy et al., 2017). Extensions to longitudinal causal data are very limited and restricted to the parametric frequentist paradigms (Robins et al., 2000; Vansteelandt and Sjolander, 2016). In particular, g-computation can be computationally intensive and intractable for dose–response estimation in the longitudinal setting with time-varying treatment and confounding and repeatedly measured outcomes, as it requires fitting sequential models of the complete data-generating process and simulating counterfactual trajectories of time-varying covariates and outcomes at a set of treatment levels. (Robins, 1986; Linero, 2022). Therefore, in this paper, we fill this literature gap by presenting a flexible yet computationally tractable generalized double non-parametric Bayesian estimation framework that estimates the entire longitudinal dose-response surface while making minimal distribution assumptions about the functional form of the continuous exposure. The Bayesian lens for longitudinal causal dose-response analysis is largely missing and remains underexplored in the literature. The Bayesian estimation framework can offer both statistical and practical advantages, such as the propagation of uncertainty through complex model structures and the ability to provide probabilistic summaries of causal effects for intuitive causal decision-making in applied research. In addition, it allows for the incorporation of subjective prior beliefs about the treatment effect and confounding from prior scientific knowledge.
@@
Our specific methodological contributions in this paper are on three fronts: i) we introduce the generalized double non-parametric Bayesian estimation framework for dose-response causal analysis with longitudinal data, integrating a similar moment-based restrictions to generalized estimating equations with Bayesian uncertainty propagation; ii) we demonstrate the validity of our approach by showing the theoretical basis behind how the approximate frequentist estimation has a natural Bayesian analogue that rests on equally weak assumptions; and iii) we provide computationally scalable algorithms and open-source code making the proposed Bayesian approach practical for large, real-world datasets for wide adoption and application beyond the motivating study.
@@
The remainder of this paper is organized as follows. In Section 2, we review methods for causal analysis of longitudinal data with repeated outcomes and introduce a nonparametric Bayesian formulation for inference when the model is specified through estimating equations. Section 3 presents simulation studies, where we compare a weighting method and an approach that includes the GPS as a covariate. In Section 4, we apply our methods to an empirical study, analyzing the causal effect of mass transit ridership on the spread of COVID-19. Finally, we conclude with some remarks in Section 5.
@@ 2 Methodology
2.1 Causal inference for longitudinal data with repeated outcomes
In causal inference for the longitudinal setting, for $t=0,\ldots,T$, let $D_{t}\in\mathbb{R}$ denote the (possibly continuous) dose at time $t$, $\bar{D}_{t}=(D_{t-c},\ldots,D_{t})$ the treatment dosing history where $1\leq c\leq t-1$ specifies the length of the relevant treatment history. Therefore, $\bar{D}_{t}$ represents the sequence of recent doses up to and including time $t$, which may influence the potential outcome $Y_{t}$. In our COVID analysis, we set $c=1$, corresponding to a first-order dosing history. However, the framework can be generalized to larger values of $c$, allowing the outcome $Y_{t}$ to depend on a longer history of past doses. $X_{t}$ a set of time-varying covariates, and $Y_{t}$ the continuous outcome measured at $t$. Using the potential outcome framework in the longitudinal setting, we set $Y_{t}(\bar{d}_{t})$ to denote the counterfactual outcome at time $t$ that would be observed if an individual were assigned the dose history $\bar{d}_{t}=(d_{1},\ldots,d_{t})$ up to time $t$. The primary focus of our inference is on average potential outcomes (APOs) for a population under a hypothetical intervention, given the dose history $\bar{d}_{T}=(d_{1},\ldots,d_{T})$ over $T$ intervals. That is, the estimand of interest is $\mathbb{E}\!\left[Y_{t}(\bar{d}_{t})\right]$. To identify the APO, we require several assumptions. First, we assume that there is no interference across units, i.e., for each individual $i$, $Y_{t,i}(\bar{d}_{1},\ldots,\bar{d}_{n})=Y_{t,i}(\bar{d}_{i})$. This is termed as the stable unit treatment value assumption (SUTVA). Second, we require that if an individual actually follows dose history $\bar{D}_{t}=\bar{d}_{t}$, then their observed outcome at each time matches the corresponding counterfactual, i.e. $Y_{t}=Y_{t}(\bar{d}_{t})$ for all $t=1,\ldots,T$, so that the observed trajectory equals the counterfactual trajectory under the realized dose sequence. Third, we assume that, at each time $t=1,\ldots,T$, conditional on the covariate $X_{t}$, past doses $\bar{D}_{t-1}$, and past outcomes $\bar{Y}_{t-1}$, the dose assignment at $t$ is independent of the future counterfactual outcomes, i.e. $Y_{t}(\bar{d}_{t})\perp\!\!\!\perp D_{t}\mid X_{t},\bar{D}_{t-1}$. Finally, for any history $\{x_{t},\bar{d}_{t-1}\}$ with positive density, we assume that the conditional density of treatment assignment satisfies $f(d_{t}\mid X_{t}=x_{t})>0$ over the range of doses considered for the target dose–response. Following the standard marginal effect parametrization of time-varying treatment and a repeatedly measured outcome, the APOs at each time $t$ is posited as
$g(\mathbb{E}\!\left[Y_{t}(\bar{d}_{t})\right])=m(\bar{d}_{t},x_{t};\,\xi),$
where $m(\cdot;\xi)$ is the conditional mean function of the treatment history $\bar{d}_{t}$ and the time-varying covariates $x_{t}$ characterized by the set of parameters $\xi$, and $g(\cdot)$ is the link function. Marginal structural models (MSMs) can be applied to directly model the marginal distribution of the outcome, while inverse probability weights are applied at the subject–visit level to construct a pseudo-population in which treatment assignment is independent of past covariates and outcome history. This reweighting removes time-dependent confounding, enabling consistent estimation of causal effects. Instead of reweighting, Graham et al. (2014) proposed a longitudinal mixed model extension of the generalized propensity score estimator to handle time-varying confounding and bi-directionality between treatment and outcome. This approach requires incorporating the generalized propensity score directly into the outcome model, $m(\bar{d}_{t},x_{t};\,\xi)$. This procedure yields consistent estimates of the APO if the propensity score model is correctly specified. Therefore, one can rewrite the MSM with functions of the estimated treatment dose densities to adjust for time-dependent confounding, i.e.,
$g(\mathbb{E}[Y_{t}\mid\bar{D}_{t},X_{t},\widehat{e}_{t}(X_{t})])=m(\bar{D}_{t},X_{t};\xi)+h\!\left(\widehat{e}_{t}(X_{t})\right),$
where $\widehat{e}_{t}(\cdot)$ denotes the estimated generalized propensity score at time $t$, and $h(\cdot)$ is a pre-specified adjustment function, which can be estimated some non-parametric method, such as splines. Therefore, the marginal treatment effect, $\xi$, can then be estimated using a generalized estimating equation (GEE) (Preisser et al., 2002), defined by the following estimating equation
@@ 2.1 Causal inference for longitudinal data with repeated outcomes
$\sum_{i=1}^{n}\mathbf{U}_{i}\left(\xi\right)={\sum\limits_{i=1}^{n}{\left({{{\partial g(\mathbb{E}[Y_{i}\mid\bar{D}_{i},X_{i},\widehat{e}_{i}(X_{i})])}\over{\partial\xi}}}\right)}^{\top}}\Sigma_{i}^{-1}\left({{Y_{i}}-g(\mathbb{E}[Y_{i}\mid\bar{D}_{i},X_{i},\widehat{e}_{i}(X_{i})])}\right)=0$ (1)
@@
where $Y_{i}$}} is a vector of observed response for ithith subject. This connects the targeting parameter $\xi$ from a link function, and $\Sigma_{i}$ is the $T\times T$ working covariance matrix for subject $i$. However, as noted in Tchetgen et al. (2012), the GEE approach consistently estimates the parameters of a correctly specified outcome model in absence of weights, irrespective of whether the working correlation structure is correct. The GEE framework offers flexibility in targeting the marginal treatment effect for repeated measurements. However, its reliance on estimating equations, rather than full likelihoods, poses challenges for conventional Bayesian analysis, as there is no explicit distributional form for the marginal mean model. This motivates the development of a fully Bayesian approach, which we outline in the next section.
2.2 Bayesian non-parametric method
@@
The Bayesian framework offers calibrated uncertainty quantification and prediction, which can render a more thorough understanding of the causal effect from a practical perspective. There is growing interest in the application of Bayesian methodology to causal problems, but most proposed methods appeal to the principles of PS-adjustment or flexible outcome modeling (Stephens et al., 2023; Li et al., 2023). An advantage of the Bayesian approach for estimation of these models is that estimated causal effects are in the form of distributions, rather than point estimates, allowing us to make probability statements about the causal quantities of interest (i.e., APOs). For example, we can specify a parametric outcome model and propensity score model and then draw inference based on the posterior sample. This is challenging in the causal setting as the GEE model (1) is only partially specified via the conditional mean without specification of the distribution model. We have to examine procedures for Bayesian inference in the case where we wish to perform analysis of an approximate model, acknowledged to be misspecified compared to the data generating model. Suppose we assume that there is a set of estimating equations $\mathbf{U}\left(\xi\right)$ of functions such that $\mathbb{E}[\mathbf{U}\left(\xi\right)|\xi]=0$, for all $\xi$. The objective is to find a non-parametric approximate model, $p(z)$, to the true data generating model $f(z|\xi)$. Thus, we can define some discrepancy $\delta(f|p)$, subject to $\int\mathbf{U}\left(\xi\right)p(z)dz=0$ and $\int p(z)dz=1$. The problem then becomes
$\min_{p}\delta(f|p)\text{ subject to }\int\mathbf{U}\left(\xi\right)p(z)dz=0,\;\;\int p(z)dz=1,\;\;\forall\xi\in\Xi.$ (2)
Suppose $p(z)$ is discrete as nonparametric probabilities as $p_{i}$. Then we can deduce the problem as the following constrained optimization problem
$\min_{p_{1},\ldots,p_{n}}\delta(f|p_{1},\ldots,p_{n})\text{ subject to }\sum_{i=1}^{n}p_{i}=1,\;\;p_{i}\geq 0,\;\;\sum_{i=1}^{n}p_{i}\mathbf{U}_{i}\left(\xi\right)=\mathbf{0}.$ (3)
From a non-parametric viewpoint, we need to consider placing the prior distribution for this non-parametric approximate model, $p$. In the next section, we link this prior specification to the Bayesian bootstrap framework and a generalized version, which enables us to derive the posterior distribution of $\xi$ by solving the constraint in (2) in a functional form.

### v3

@@ Abstract
Existing causal methods for time-varying exposure and time-varying confounding focus on estimating the average causal effect of a time-varying binary treatment on an end-of-study outcome, offering limited tools for characterizing marginal causal dose–response relationships under continuous exposures. We propose a scalable, nonparametric Bayesian framework for estimating marginal longitudinal causal dose–response functions with repeated outcome measurements. Our approach targets the average potential outcome at any fixed dose level and accommodates time-varying confounding through the generalized propensity score. The proposed approach embeds a Dirichlet process specification within a generalized estimating equations structure, capturing temporal correlation while making minimal assumptions about the functional form of the continuous exposure. We apply the proposed methods to monthly metro ridership and COVID-19 case data from major international cities, identifying causal relationships and the dose–response patterns between higher ridership and increased case counts.
Keywords: Dose-response relationship; Bayesian bootstrap; Dirichlet processes; Longitudinal data analysis; COVID-19; Transportation engineering
@@ 1 Introduction
Using the standard causal inference notation, we define our problem in relation to the triple $(Y,D,X)$, indexed by $t$ to denote time: where $Y$ is the outcome (COVID-19 cases in the city), $D$ is the treatment (urban mass transit ridership), $X$ is a set of time-varying confounders and $U$ is a set of confounders that are time-invariant. We are interested in the causal effects of $D$ on $Y$. Therefore, we face two key challenges in estimating this marginal causal effect. First, we must adjust for the existence of confounding influences, $X$ and $U$, which determine both transit ridership and disease propagation. These include time-varying (i.e., $X$) factors such as the severity of containment measures on human mobility (e.g. lockdown measures), urban mobility characteristics, and vaccination rates, along with a set of unmeasured time-invariant (i.e. $U$) city level attributes such as characteristics of the built-environment, urban culture, and baseline socio-demographics. Second, there are dynamic effects presented such that past outcomes act as contemporaneous confounders. Thus, we observe a repeatedly measured outcome, a time-dependent treatment (dose) and a set of dynamic time-varying confounders.
@@
Most existing methods for causal estimation of the effect of a continuous-exposure are developed for the point-treatment framework, including the generalized propensity score (GPS)-based approaches (Robins et al., 2000; Hirano and Imbens, 2004; Moodie and Stephens, 2012; Galvao and Wang, 2015), g-computation and outcome regression-based approaches (Hill, 2011), and a doubly robust estimator considering combining both treatment and outcome models (Kennedy et al., 2017). Extensions to longitudinal causal data are very limited and restricted to the parametric frequentist paradigms (Robins et al., 2000; Vansteelandt and Sjolander, 2016). In particular, g-computation can be computationally intensive and intractable when estimating treatment regime effects in longitudinal settings with high-dimensional time-varying confounding, as it requires fitting sequential models of the complete data-generating process and simulating counterfactual trajectories of time-varying covariates and outcomes at a set of treatment levels (Robins, 1986; Linero, 2022). On the other hand, the Bayesian lens for longitudinal causal dose-response analysis is largely missing and remains underexplored in the literature. The Bayesian estimation framework can offer both statistical and practical advantages, including the propagation of uncertainty through complex model structures and the ability to provide probabilistic summaries of causal effects for intuitive causal decision-making in applied research, which facilitates decision-making in applied settings such as our ridership–COVID study. In addition, frequentist methods such as generalized estimating equations (GEEs) do not naturally accommodate prior information. These features motivate a Bayesian alternative that retains the scalability of estimating-equation methods while enabling posterior summaries of longitudinal dose–response effects. The incorporation of prior beliefs under the proposed Bayesian approach can be particularly valuable in biomedical and policy applications where historical evidence or expert knowledge is available. Moreover, unlike approaches that rely on large-sample approximations, Bayesian inference provides coherent uncertainty quantification even in small-sample settings, which commonly arise in dose–response analyses. These considerations motivate us to develop Bayesian alternatives for longitudinal causal dose–response modeling. In this paper, therefore, we fill this literature gap by introducing a flexible yet computationally tractable non-parametric Bayesian estimation framework that estimates the entire longitudinal dose-response surface while making minimal distribution assumptions about the functional form of the continuous exposure.
@@
Our specific methodological contributions in this paper are on three fronts: i) we propose a nonparametric Bayesian framework using the Dirichlet process (DP) for marginal causal dose–response estimation with longitudinal data, which integrates GEEs with coherent Bayesian uncertainty quantification; ii) we demonstrate how the proposed DP–based framework accommodates multiple causal adjustment strategies, including covariate-adjusted outcome regression and inverse probability weighted marginal structural models; and iii) we provide computationally scalable algorithms and open-source code (https://github.com/yumcgill/Bayes_DoseResponse) making the proposed Bayesian approach practical for large, real-world datasets for wide adoption and application beyond the motivating study.
@@
The remainder of this paper is organized as follows. In Section 2, we review methods for causal analysis of longitudinal data with repeated outcomes and introduce a non-parametric Bayesian formulation for inference when the model is specified through GEEs. Section 3 presents simulation studies, comparing a weighting-based method with an approach that incorporates the GPS as a covariate. In Section 4, we apply our methods to a real data example, analyzing the causal effect of mass transit ridership on the spread of COVID-19. Finally, we conclude with some remarks in Section 5.
@@ 2 Methodology
2.1 Causal inference for marginal dose–response functions with repeated outcomes
In causal inference for the longitudinal setting, for $t=1,\ldots,T$, let $D_{t}\in\mathbb{R}$ denote the (possibly continuous) dose at time $t$, Let $X_{t}$ denote a set of time-varying observed covariates, and let $U$ denote an unobserved time-invariant baseline covariate that may affect both treatment assignment and outcomes. Therefore, the observed data consist of the longitudinal process $Z=\{(Y_{t},X_{t},D_{t}):t=1,\ldots,T\}$. Using the potential outcome framework for longitudinal data, we define $Y_{t}(d_{t})$ as the counterfactual outcome at time $t$ that would be observed if a study unit were exposed to the dose $d_{t}$ at time $t$. Our primary estimand is a marginal dose–response function, defined as
$\mu(d)=\mathbb{E}\!\left\{Y(d)\right\},$
where $Y(d)$ denotes the potential outcome that would be observed if, at each time point, the exposure were set to the constant level $d$. This estimand represents the average potential outcome (APO) under an exposure level $d$. In other words, the estimand depends only on the marginal level of exposure $d$, rather than on the full longitudinal treatment path used to achieve it. This marginal dose–response parameterization was commonly used in transportation engineering. As in the transportation setting, it is neither feasible nor substantively meaningful to define a single, well-specified longitudinal treatment regime. Instead, our scientific goal is to quantify the APO at a given ridership on COVID-related cases. Therefore, we believe it is suitable to estimate the dose-response relationship in our COVID analysis. For example it was employed by Graham et al. (2014) in their analysis to quantify the effect of the network capacity expansions on urban traffic volumes. Therefore, we adopt this estimand to characterize the dose–response relationship in our COVID analysis. In Appendix B, an extension of this framework is provided for estimating regime-specific APO associated with identification results.
To identify the APO $\mu(d)=\mathbb{E}\{Y(d)\}$, we invoke the following assumptions. First, we assume no interference across units, i.e., for each study unit $i$ ($i=1,\ldots,n$), $Y_{i}(d_{1},\ldots,d_{n})=Y_{i}(d_{i})$, corresponding to the stable unit treatment value assumption (SUTVA). Second, we assume consistency, that is, if an study unit is observed to receive dose $D_{t}=d$ at time $t$, then the observed outcome satisfies $Y_{t}=Y_{t}(d)$. Third, we assume marginal ignorability of treatment assignment: at each time $t=1,\ldots,T$, $Y(d)\;\perp\!\!\!\perp\;D_{t}\mid X_{t}$. This assumption does not condition on past dose, and allows past dose to affect both $D$ and $Y$ indirectly, provided those effects are captured by $X$. Fourth, we assume positivity. For any covariate value $x_{t}$ with positive density, $f(d_{t}\mid X_{t}=x_{t})>0$ over the range of doses considered. Finally, we assume that unobserved time-invariant heterogeneity $U$ is mean-independent of the observed time-varying covariates, i.e., $\mathbb{E}(U\mid X_{t})=0$. Figure 1 illustrates a general longitudinal causal structure consistent with these assumptions, allowing for treatment–confounder feedback and unobserved time-invariant heterogeneity and targeting a marginal estimand.
$X_{t}$$D_{t}$$Y_{t}$$U$
Figure 1: Directed acyclic graph for identification of the marginal dose–response function $\mu(d)=\mathbb{E}\{Y(d)\}$ under marginal ignorability. Observed covariates $X_{t}$ block all backdoor paths between exposure $D_{t}$ and outcome $Y_{t}$, while unobserved time-invariant heterogeneity $U$ is orthogonal to $X_{t}$.
Following this marginal effect parameterization, the APO can be modeled as
$g\!\left(\mathbb{E}\!\left\{Y(d)\right\}\right)=m(d;\xi),$
where $g(\cdot)$ is a known link function and $m(\cdot;\xi)$ is a marginal dose–response model indexed by parameters $\xi$. Marginal structural models (MSMs) estimate $\mu(d)$ by reweighting observations using inverse probability weights to remove time-dependent confounding. As an alternative to reweighting, Graham et al. (2014) proposed incorporating the GPS directly into the outcome model. Under correct specification of the GPS model, this approach yields consistent estimation of the marginal dose-response function. Accordingly, the outcome model may be written as
$g\!\left(\mathbb{E}[Y_{t}\mid D_{t},X_{t},\widehat{e}_{t}(X_{t})]\right)=m(D_{t};\xi)+h\!\left(\widehat{e}_{t}(X_{t})\right),$
where $\widehat{e}_{t}(\cdot)$ denotes the estimated GPS at time $t$ and $h(\cdot)$ is a prespecified adjustment function, which may be estimated flexibly (e.g., using splines). By integrating over time and covariate, estimation of $\xi$ yields the time-invariant APO $\mu(d)$. For a continuous exposure, the GPS is defined as $e_{t}(X_{t})=f(D_{t}\mid X_{t})$, where $f(D_{t}\mid X_{t})$ denotes the conditional density of the treatment given covariates. In practice, the GPS is estimated by specifying a marginal mean model for $D_{t}$ given $X_{t}$ and estimating its parameters using the GEE to account for within-subject dependence across repeated treatment measurements. Importantly, although the treatment assignment may depend on an unobserved time-invariant factor $U$, the GEE estimator targets the marginal mean $E(D_{t}\mid X_{t})$, which integrates over the distribution of $U$. Under the assumption that $E(U\mid X_{t})=0$, the resulting estimating equations are unbiased, and the GEE estimator yields consistent estimates of the marginal treatment model parameters regardless of the working correlation structure. The estimated marginal mean and variance are then used to evaluate the conditional density of $D_{t}$ (e.g., assuming normality) at the observed exposure level, yielding a consistent estimate of the GPS. Incorporating this consistently estimated GPS into the outcome GEE therefore yields consistent estimation of the APO. Appendix A provides formal results demonstrating that unobserved time-invariant effects $U$ are integrated out in the marginal treatment model and therefore do not bias the GEE-based estimation of the GPS. Under assumptions, the dose–response function $\mu(d)$ is identified by
$\mathbb{E}[Y_{t}(d)]=\mathbb{E}_{X_{t}}\left[\mathbb{E}\left[Y_{t}(d)\right]\right]=\mathbb{E}_{X_{t}}\left[\mathbb{E}\left[Y_{t}\mid D_{t}=d,e_{t}(d,X_{t})\right]\right],\qquad d\in\mathcal{D},$ (1)
where $e_{t}(d,X_{t})=f(D_{t}=d\mid X_{t})$ and can be consistently estimated via the GEE treatment model. Therefore, the marginal treatment effect associated with parameter $\xi$, can then be estimated using a GEE (Preisser et al., 2002), defined by the following estimating equation
@@ 2.1 Causal inference for marginal dose–response functions with repeated outcomes
$\sum_{i=1}^{n}\mathbf{U}_{i}(\xi)\equiv\sum_{i=1}^{n}\left(\frac{\partial\mu_{i}(\xi)}{\partial\xi}\right)^{\top}\Sigma_{i}^{-1}\left(Y_{i}-\mu_{i}(\xi)\right)=0,\;\;\;\mu_{i}(\xi)=\mathbb{E}[Y_{t}\mid D_{t},X_{t},\widehat{e}_{t}(X_{t})]$ (2)
@@
where $Y_{i}$}} is a vector of observed response for ithith study unit. This connects the targeting parameter $\xi$ from a link function, and $\Sigma_{i}$ is the $T\times T$ working covariance matrix for study unit $i$. However, as noted in Tchetgen et al. (2012), the GEE approach consistently estimates the parameters of a correctly specified outcome model in absence of weights, irrespective of whether the working correlation structure is correct. The GEE framework offers flexibility in targeting the marginal treatment effect for repeated measurements. However, its reliance on estimating equations, rather than full likelihoods, poses challenges for conventional Bayesian analysis, as there is no explicit distributional form for the marginal mean model. This motivates the development of a fully Bayesian approach, which we outline in the next section.
2.2 Bayesian non-parametric method for GEE-based causal estimands
@@
The Bayesian framework offers calibrated uncertainty quantification and prediction, which can render a more thorough understanding of the causal effect from a practical perspective. There is growing interest in the application of Bayesian methodology to causal problems, but most proposed methods appeal to the principles of PS-adjustment or flexible outcome modeling (Stephens et al., 2023; Li et al., 2023). An advantage of the Bayesian approach for estimation of these models is that estimated causal effects are in the form of distributions, rather than point estimates, allowing us to make probability statements about the causal quantities of interest (i.e., APOs). In fully parametric Bayesian analyses, this is typically achieved by specifying likelihood models for both the outcome and treatment assignment mechanisms. However, such an approach is not directly applicable in the present setting, as the GEE in (2) specifies only the conditional mean structure of the outcome and does not correspond to a fully specified likelihood. Consequently, Bayesian inference in this context must proceed under an approximate (or partially specified) model, which may be misspecified relative to the true data-generating mechanism. Rather than modeling the full joint distribution of the data, inference targets a causal estimand defined through unbiased estimating equations. This perspective aligns with the generalized Bayesian framework and recent work on Bayesian inference for loss functions (Luo et al., 2023). We define the target causal parameter $\xi_{0}$ at the population level through the moment condition
$\mathbb{E}_{F_{0}}\!\left\{\mathbf{U}(\xi_{0})\right\}=0,$ (3)
where $F_{0}$ denotes the true, unknown data-generating distribution and $\mathbf{U}(\cdot)$ is the GEE function defined in (2). Equivalently, $\xi_{0}$ may be characterized as the minimizer of an expected loss function,
$\xi_{0}=\arg\min_{\xi\in\Xi}\;\mathbb{E}_{F_{0}}\!\left\{\ell(Z;\xi)\right\},$ (4)
where $\ell(\cdot)$ is a differentiable loss function whose gradient yields the estimating equations in (3). This loss-based definition is consistent with the generalized Bayesian framework of Luo et al. (2023) and provides a precise definition of the causal estimand without requiring a likelihood. In the Bayesian non-parametric framework, uncertainty about $F_{0}$ is represented through a posterior distribution over probability measures. If we assume that the data points are realizations from a multinomial model on the finite set of the observed data $\left\{z_{1},\ldots,z_{n}\right\}$ with unknown probability $\mathbf{p}=\left(p_{1},\ldots,p_{n}\right)$, and assuming a priori that $\mathbf{p}\sim\text{Dirichlet}\left(\alpha,\ldots,\alpha\right)$, then, a posteriori $\mathbf{p}\sim\text{Dirichlet}\left(\alpha+1,\ldots,\alpha+1\right)$. The standard Bayesian bootstrap is obtained under the improper specification $\alpha=0$ (Rubin, 1981). Therefore, if we repeatedly sample $\mathbf{p}\sim\text{Dirichlet}\left(1,\ldots,1\right)$, then
$F^{(b)}=\sum_{i=1}^{n}p_{i}^{(b)}\delta_{Z_{i}}.$
For each posterior draw $F^{(b)}$, the corresponding parameter draw $\xi^{(b)}$ is defined as the solution to the posterior predictive estimating equation
$\int\mathbf{U}(\xi^{(b)})\,dF^{(b)}(z)=\sum_{i=1}^{n}p_{i}^{(b)}\mathbf{U}_{i}(\xi^{(b)})=0.$ (5)
This construction induces a posterior distribution over $\xi$ through the posterior over $F$, yielding valid Bayesian uncertainty quantification for GEE-defined causal estimands (Newton and Raftery, 1994; Chamberlain and Imbens, 2003). Importantly, the definition of $\xi_{0}$ in (3) and the Bayesian bootstrap implementation in (5) represent equivalent formulations of the same inferential target. The former defines the causal estimand at the population level, while the latter provides a practical posterior sampling scheme. This approach yields a coherent Bayesian non-parametric framework for inference on GEE-based causal parameters, accommodates model misspecification, and avoids the need to specify a full likelihood while retaining the interpretability of marginal causal estimands.